# Match, Compare, or Select? An Investigation of Large Language Models for Entity Matching

### v2

@@ Appendix B Detailed Results of Open-Source LLMs under Different Strategies
Mistral-Instruct-8x7B Matching 77.67 34.76 67.20 60.09 82.26 53.57 72.99 50.57 62.39
@@
Llama-3-8B Matching 31.01 21.97 19.27 19.27 44.78 40.24 31.55 23.91 29.00
Comparing 80.06 61.27 84.85 72.54 80.13 76.36 79.82 80.29 76.91
Selecting 74.37 49.50 78.91 68.79 76.27 54.77 69.66 42.33 64.33
Qwen-2-7B Matching 63.41 47.33 68.35 52.46 82.89 55.54 71.84 55.06 62.11
@@
Table 5: F1 score of open-source LLMs under different strategies.
Appendix C Ablation Study on Each Dataset
Figure 8 shows the performance of ComEM with respect to the varying number of top candidates retained for further selection. Similar to Figure 7, recall increases and precision decreases as $k$ increases. For the simplest dataset “DBLP-ACM”, F1 achieves the highest value at $k=1$. For some other datasets, F1 changes dramatically as $k$ goes from 2 to 5. How to tune $k$ to balance the LLM capabilities and the actual situation is a direction that could be explored. One possible solution might be to dynamically adjust the number according to the similarity scores of potential matches.

### v3

@@ Appendix C Detailed Results of Open-Source LLMs under Different Strategies
ComEM 70.33 61.52 83.12 78.05 87.10 76.85 83.54 66.97 75.94
Mixtral-Instruct-8x7B Matching 77.67 34.76 67.20 60.09 82.26 53.57 72.99 50.57 62.39
@@
ComEM 78.34 69.24 84.29 80.97 85.80 76.81 78.39 78.44 79.03
Llama-3.1-8B-Instruct Matching 53.97 29.93 22.97 26.54 80.14 62.00 65.94 40.56 47.75
Comparing 84.41 55.27 85.80 75.49 72.20 70.33 78.39 85.30 75.90
Selecting 78.29 68.00 81.16 75.57 79.65 75.68 77.45 74.82 76.33
ComEM 80.86 70.86 84.57 81.14 85.01 79.34 79.53 80.06 80.17
Qwen2-7B-Instruct Matching 63.41 47.33 68.35 52.46 82.89 55.54 71.84 55.06 62.11
@@
ComEM 82.46 70.69 86.68 82.68 88.26 79.22 80.06 79.88 81.24
Table 6: F1 score of open-source LLMs under different strategies and ComEM.